OpenStudy (anonymous):

sinA, cosA, tanA are in GP, the the value of cot^6(A) - cot^2(A)

hartnn (hartnn):

i am getting 1

hartnn (hartnn):

did u reach to the point 1/cos^3 A - 1/ cos A ??

hartnn (hartnn):

how much of it did u solve ? and is the answer = 1 ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

can u say d whole procedure

hartnn (hartnn):

i can, but i need to know how much did u solve, then i'll continue from there

hartnn (hartnn):

did u get sin^2 A = cos ^3 A ?

OpenStudy (anonymous):

no

hartnn (hartnn):

u know the meaning of GP ?

OpenStudy (anonymous):

yeah... of course

hartnn (hartnn):

so sin A tan A = cos^2 A can u get sin^2 A = cos ^3 A from here ?

OpenStudy (anonymous):

ya... i got

hartnn (hartnn):

now write cot A in terms of cosA and sin A

OpenStudy (anonymous):

cos A/ sin A = cot A

OpenStudy (anonymous):

COT^2 A = COS A

hartnn (hartnn):

i mean write it here :cot^6(A) - cot^2(A) cos^6 A/ sin^6 A - cot^2 A / sin^2 A

hartnn (hartnn):

cos^6 A/ sin^6 A - cos^2 A / sin^2 A

OpenStudy (anonymous):

i did do it

hartnn (hartnn):

so sin^6 A = cos^9 A right ?

OpenStudy (anonymous):